\begin{thebibliography}{99}

\bibitem{AH}

L. Alvarez-Gaum\'e and S. F. Hassan, Introduction to $S$-Duality in $N = 2$ Supersymmetric Gauge Theories, arXiv:hep-th/9701069.


\bibitem{Bilal}

A. Bilal, DUALITY IN $N = 2$ SUSY $SU(2)$ YANG-MILLS THEORY: A pedagogical introduction to the work of Seiberg and Witten. arXiv:hep-th/9601007.

\bibitem{Borcherds}

R. E. Borcherds, Automorphic forms with singularities on Grassmannians, Invent math 132, 491–562 (1998). \url{https://doi.org/10.1007/s002220050232}.


\bibitem{Zagier}
J. H. Bruinier, G. van der Geer, G. Harder and D. Zagier, The 1--2--3 of Modular Forms, Lectures at a Summer School in Nordfjordeid, Norway. Springer.


\bibitem{CarlsonMullerPeters}
J. Carlson, S. M\"uller-Stach, and C. Peters. Period Mappings and Period Domains. Cambridge University Press.

\bibitem{Chandra}
K. Chandrasekharan, Elliptic Functions, Grundlehren der mathematischen Wissenschaften, 281, Springer-Verlag.

\bibitem{Diamond}

F. Diamond and J. Shurman, A First Course in Modular Forms, GTM, vol. 228, Springer (2005).

\bibitem{MarkGross}
M. Gross, D. Huybrechts and D. Joyce, Calabi-Yau Manifolds and Related Geometries. Springer.

\bibitem{Hille}

E. Hille, Ordinary differential equations in the complex domain, Dover Publications (1997).


\bibitem{KimYang}
M. Kim and W. Yang, Mirror symmetry, mixed motives and $\zeta(3)$. arXiv:1710.02344.


\bibitem{SS}

A. Sebbar and H. Saber, Automorphic Schwarzian equations. Forum Mathematicum | Volume 32: Issue 6. DOI: \url{https://doi.org/10.1515/forum-2020-0025}. arXiv: 2002.00493.


\bibitem{SeibergWitten}

N. Seiberg, E. Witten, Electric-magnetic duality, monopole condensation, and confinement in N=2 supersymmetric Yang-Mills theory. Nucl.Phys. B426 (1994), 19-52. arXiv:  hep-th/9407087.


\bibitem{ZagierAT}

D. Zagier, Arithmetic and topology of differential equations, in Proceedings of the Seventh European
Congress of Mathematics (Berlin, July 18–22, 2016), Editors V. Mehrmann, M. Skutella, European Mathematical Society, Berlin, 2018, 717–776.

\bibitem{OEIS}

A035016,  The On-Line Encyclopedia of Integer Sequences. \url{https://oeis.org/A035016}.



\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{amssymb,amsthm,amsmath,amstext,amsxtra}
\usepackage{mathrsfs,bm}
\usepackage{fullpage}
\usepackage[all]{xy}
\usepackage{booktabs}
\usepackage{mathtools}
\usepackage{hyperref}
\hypersetup{colorlinks=true,urlcolor=blue,citecolor=blue,linkcolor=blue}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{comment}
\usepackage{multirow}
\usepackage{colonequals}
\usepackage{tikz}
\usepackage{marginnote}
\usepackage[export]{adjustbox}

\usepackage{tikz}
\usepackage{tikz-cd}
\usepackage{xcolor}

\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\begin{document}

\title{Seiberg-Witten theory and modular lambda function}

\author{Wenzhe Yang}
\address{SITP, Physics Department, Stanford University, Stanford, CA, 94305}
\email{yangwz@stanford.edu}

\begin{abstract}
In this paper, we will apply the tools from number theory and modular forms to the study of the Seiberg-Witten theory. We will express the holomorphic functions $a, a_D$, which generate the lattice $Z=n_e a+n_m a_D, (n_e, n_m) \in \mathbb{Z}^2$ of central charges, in terms of the periods of the Legendre family of elliptic curves. Thus we will be able to compute the transformations of the quotient $a_D/a$ under the action of the modular group $\text{PSL}(2,\mathbb{Z})$. We will show the Schwarzian derivative of the quotient $a_D/a$ with respect to the complexified coupling constant is given by the theta functions. We will also compute the scalar curvature of the moduli space of the $N=2$ supersymmetric Yang-Mills theory, which is shown to be asymptotically flat near the perturbative limit.

\end{abstract}

\maketitle
\setcounter{tocdepth}{1}
\vspace{-13pt}
\vspace{-13pt}

\vspace*{0.2in}

\tableofcontents
\section{Introduction}

In the fundamental paper \cite{SeibergWitten} by Seiberg and Witten, the low-energy effective action of the $N=2$ supersymmetric Yang-Mills theory with gauge group $\text{SU}(2)$ has been determined by using the holomorphic properties of the theory together with their monodromy behaviors near certain singularities. Their beautiful work is now called the Seiberg-Witten theory. In this paper, we will apply the tools from number theory, e.g. modular lambda function, Schwarzian derivative and Great Picard's Theorem, to study the Seiberg-Witten theory, which provide new insights into the theory.

First, let us briefly introduce some well-known results of Seiberg-Witten theory that will be needed in this paper. This section is never meant to be a review of Seiberg-Witten theory, not even a brief one. So we will only introduce the results, while leave all the details to the original paper \cite{SeibergWitten}. The readers who are not familiar with the $N=2$ supersymmetric Yang-Mills theory can also consult the review papers \cite{AH, Bilal}.

 The $N=2$ SUSY multiplet contains a massless gauge field $A_\mu$, two Weyl fermions $\lambda$ and $\psi$, and a scalar field $\phi$, all of which live in the adjoint representation of the gauge group $\text{SU}(2)$. The classical potential of the pure $N=2$ theory is given by
\begin{equation}
V(\phi)=\frac{1}{2} \text{Tr}[\phi, \phi^\dagger].
\end{equation}
Unbroken SUSY will require that we have 
\begin{equation}
V(\phi) =0
\end{equation}
in the vacuum. But this does not mean that $\phi$ itself is zero, since it is sufficient that if $\phi$ and $\phi^\dagger$ commute. 

Up to a gauge transformation, the scalar field $\phi$ can be taken to be
\begin{equation}
\phi =\frac{1}{2} a \sigma_3,
\end{equation}
where $\sigma_3$ is the third Pauli matrix. For a given vacuum, $a$ must be a constant. But a gauge transformation from the Weyl group of $\text{SU}(2)$ can still transform $a$ to $-a$, so they are gauge equivalent. Hence the gauge invariant quantity that parametrizes the inequivalent vacua is
\begin{equation}
u = \frac{1}{2}a^2=\text{Tr} \,\phi^2.
\end{equation}
The space of gauge inequivalent vacua is also called the moduli space $\mathscr{M}$ of the $N=2$ theory, which has $u$ as its coordinate. Naively, one can consider $\mathscr{M}$ as the complex plane $\mathbb{C}$. However, this moduli space $\mathscr{M}$ has special singularities, and the behaviors of the $N=2$ theory near these singularities will determine the low-energy effective action \cite{SeibergWitten}.

More explicitly, the low energy effective Lagrangian of the $N=2$ supersymmetric Yang-Mills theory is completely determined by a prepotential $\mathcal{F}(a)$, which is a multi-valued holomorphic function. The effective complexified coupling constant of the $N=2$ theory is given by
\begin{equation}
\tau_{sw}(a) = \frac{\partial^2 \mathcal{F}}{\partial a^2}.
\end{equation} 
While a natural choice of metric for the moduli space $\mathscr{M}$ is 
\begin{equation} \label{eq:metriconModuliSpace}
(ds)^2=\text{Im} \,\tau_{sw}(a) \, da d \overline{a},
\end{equation}
which works at least in a small neighborhood of the perturbative limit $u=\infty$. The dual variable $a_D$ is given by
\begin{equation}
a_D= \partial \mathcal{F} /\partial a,
\end{equation}
and the complexified coupling constant satisfies
\begin{equation}
\tau_{sw}(a) =\frac{d a_D}{d a}.
\end{equation}
In particular, the metric $(ds)^2$ on the moduli space can also be expressed as 
\begin{equation} \label{eq:metriconmodulispace}
(ds)^2 =\text{Im} \,da_D d\overline{a} =-\frac{i}{2} \left( da_D d\overline{a} -da d\overline{a}_D \right).
\end{equation}

The key insight of Seiberg and Witten is that the multi-valued holomorphic functions $a(u)$ and $a_D(u)$ can be realized as the period integrals of a meromorphic oneform defined for a family of elliptic curves \cite{SeibergWitten}. More explicitly, let $\mathcal{E}_u$ be the family of elliptic curves defined by the cubic equation
\begin{equation} \label{eq:swfamilyelliptic}
\mathcal{E}_u:y^2=(x-1)(x+1)(x-u).
\end{equation}
On every elliptic curve $\mathcal{E}_u$, there is a meromorphic oneform $\Omega^{sw}_u$, called the Seiberg-Witten oneform, that is defined by
\begin{equation} \label{eq:swoneformU}
\Omega^{sw}_u=\frac{\sqrt{2}}{2 \pi} \left( \frac{x dx}{y} -\frac{u dx}{y} \right).
\end{equation}
The residue of $\Omega^{sw}_u$ at its pole is zero, therefore it defines an element of the cohomology group $H^1(\mathcal{E}_u,\mathbb{C})$ \cite{SeibergWitten}. The underlying differentiable manifold of $\mathcal{E}_u$ is the torus. From the paper \cite{SeibergWitten}, there exists a basis $\{ \gamma^{sw}_0, \gamma^{sw}_1 \}$ for the homology group $H_1(\mathcal{E}_u,\mathbb{Z})$ such that 
\begin{equation} \label{eq:aaDperiods}
a(u)=\int_{\gamma^{sw}_0} \Omega^{sw}_u,~a_D(u)=\int_{\gamma^{sw}_1}\,\Omega^{sw}_u.
\end{equation}
In fact, the mass-lattice of the BPS states of the $N=2$ theory is given by
\begin{equation} \label{eq:BPSmasslattice}
\{n_e a+n_m a_D:(n_e,n_m) \in \mathbb{Z}^2 \}.
\end{equation}
Furthermore, the complexified coupling constant $\tau_{sw}(u)$ is just the period of the elliptic curve $\mathcal{E}_u$ that is given by
\begin{equation} \label{eq:CCCu}
\tau_{sw}(u) =\frac{d a_D/du}{d a/du}.
\end{equation}
The quotient $a_D/a$ plays an important role in the Seiberg-Witten theory, especially in the study of the wall-crossing of BPS states \cite{AH, Bilal, SeibergWitten}.

The motivation of this paper is that since the transformation
\begin{equation} \label{eq:ellTransformation}
x \mapsto 2x-1,~y\mapsto 2 \sqrt{2}\,y,~u \mapsto 2 \lambda-1,
\end{equation}
sends the family \eqref{eq:swfamilyelliptic} to the Legendre family of elliptic curves
\begin{equation}
\mathscr{E}_{\lambda}: y^2=x(x-1)(x-\lambda),
\end{equation}
could the tools in number theory provide any new insights into the Seiberg-Witten theory?

We will see the answer is yes! The most significant result we have obtained through this approach is that the Schwarzian derivative of the quotient $a_D/a$ with respect to the complexified coupling constant $\tau_{sw}$ is given by the theta functions
\begin{equation} \label{eq:swperiodSchwarzian}
\{a_D/a,  \tau_{sw}\} =\frac{\pi^2}{2} \theta^8_4(0,e^{\pi i \tau_{sw}}).
\end{equation}
Recall that the Schwarzian derivative of a holomorphic function $f(z)$ with respect to the variable $z$ is defined by
\begin{equation}
\{f, z \}=\left( \frac{f''(z)}{f'(z)} \right)' -\frac{1}{2}\left( \frac{f''(z}{f'(z)} \right)^2 =\frac{f'''(z)}{f'(z)} -\frac{3}{2}\left( \frac{f''(z)}{f'(z)} \right)^2.
\end{equation}
The Schwarzian derivative has many important properties \cite{Hille, SS}, e.g.
\begin{enumerate}

\item It is invariant under the M\"obius transformation
\begin{equation} \label{eq:SDMobiusTransformation}
\left\{\frac{p f+q}{r f +s}, z \right\} =\{f,  z\},\,\,\,\,
\left(
\begin{array}{cc}
 p & q \\
 r & s \\
\end{array}
\right) \in \text{GL}(2,\mathbb{C}).
\end{equation}

\item If $w$ is a function of $z$, then we have
\begin{equation} \label{eq:SDcocycle}
\{f,z\} =\{f, w \} \left( \frac{dw}{dz} \right)^2 + \{w, z\}.
\end{equation}

\item If $w$ is related to $z$ by a M\"obius transformation
\begin{equation}
w=\frac{pz+q}{rz+s},~
\left(
\begin{array}{cc}
 p & q \\
 r & s \\
\end{array}
\right) \in \text{GL}(2,\mathbb{C}),
\end{equation}
then we have
\begin{equation}\label{eq:SDvariabletransformation2}
\{f,z\} =\{ f,w\} \frac{(ps-qr)^2}{(rz+s)^4}.
\end{equation}

\item Given two holomorphic functions $f(z)$ and $g(z)$, we have
\begin{equation} \label{eq:SDvariableTransformation3}
\{ f, z\}=\{ g,z\} \iff f(z)=\frac{p g(z)+q}{r g(z) +s}\,\,\,\text{for some }
\left(
\begin{array}{cc}
 p & q \\
 r & s \\
\end{array}
\right) \in \text{GL}(2,\mathbb{C}).
\end{equation}
\end{enumerate}
These properties of the Schwarzian derivative immediately tell us that the equation \eqref{eq:swperiodSchwarzian} is an essential equation that characterize the central charge lattice \eqref{eq:BPSmasslattice}, and more importantly the Seiberg-Witten theory. This equation prompts us to ask whether there exists an automorphic representation that can be constructed from the the quotient $a_D/a$ which is related to the counting of the BPS states of the $N=2$ supersymmetric Yang-Mills theory. It is also very interesting to find out whether the equation \eqref{eq:swperiodSchwarzian} admits any physical interpretations.

In this paper, we will also apply the Great Picard's Theorem to study the behaviors of the quotient $a_D/a$ near the non-perturbative limit $u= \pm 1$, based on which we prove that the quotient $a_D/a$ indeed can take real values. Even though this is already a well-known property, but a mathematical proof could at least serve as an assurance. We will also compute the scalar curvature of the metric \eqref{eq:metriconModuliSpace} on the moduli space $\mathscr{M}$. We will show that near the perturbative limit $u=\infty$, the scalar curvature approaches 0 exponentially, while near the non-perturbative limits $u=\pm 1$, the scalar curvature approaches infinity exponentially. Therefore near the perturbative limit $u=\infty$, the moduli space $\mathscr{M}$ is asymptotically flat.

The outline of this paper is as follows. In Section \ref{sec:legendrefamily}, we will review some well-known results of the Legendre family of elliptic curves, e.g. its periods and the Picard-Fuchs equation. In Section \ref{sec:modularlambdafunction}, we will discuss the properties of the modular lambda function and the theta functions. In Section \ref{sec:seibergwittenperiods}, we will look at the meromorphic Seiberg-Witten oneform and its periods. We will find out their relations to the periods of the Legendre family of elliptic curves. In Section \ref{sec:SchwarzianDerivative}, we will study the transformations of the periods of the meromorphic Seiberg-Witten oneform under the action of the modular group $\text{PSL}(2,\mathbb{Z})$. We will show the Schwarzian derivative of the quotient $a_D/a$ with respect to the complexified coupling constant $\tau_{sw}$ is given by the theta functions. In Section \ref{sec:greatpicards}, we will apply the Great Picard's Theorem to study the behaviors of the quotient $a_D/a$ near the non-perturbative limit $u=\pm 1$. In Section \ref{sec:metric}, we will compute the scalar curvature for the moduli space $\mathscr{M}$ of the $N=2$ theory. In Section \ref{sec:conclusion}, we will summarize the results of this paper and raise several interesting open questions.

\section{The Legendre family of elliptic curves} \label{sec:legendrefamily}

In this section, we will review some well-known results about the Legendre family of elliptic curves, which are very elementary. They are included here solely to make this paper as self-contained as possible.

The Legendre family of elliptic curves is defined by the cubic equation
\begin{equation} \label{eq:legendrefamilyequation}
\mathscr{E}_{\lambda}: y^2=x(x-1)(x-\lambda),
\end{equation}
where $\lambda$ is a free parameter. The singular fibers of this family are over the points $\lambda=0,1,\infty$. When $\lambda$ is 0 or 1, this family degenerates to a nodal cubic. Let us now briefly recall how to construct a complex curve from the equation \eqref{eq:legendrefamilyequation} using branch cuts  \cite{CarlsonMullerPeters}. First, cut the complex plane $\mathbb{C}$ along two lines: 0 to $\lambda$ and 1 to $\infty$. Second, take a second copy of the complex plane $\mathbb{C}$ and cut it along the same lines. Then, glue the two copies of the complex planes along the corresponding branch cuts. The result is a smooth torus, which has a complex structure induced by the cubic equation \eqref{eq:legendrefamilyequation}.

On every smooth elliptic curve $\mathscr{E}_\lambda$, there exists a nowhere vanishing oneform $\Omega_\lambda$ which on the open locus where $y \neq 0$ is defined by
\begin{equation} \label{eq:LengendreOneform}
\Omega_\lambda=\frac{dx}{2\pi y}.
\end{equation}
It is a straightforward exercise to check that $dx/(2\pi y)$ actually extends to a nowhere vanishing oneform on $\mathscr{E}_{\lambda}$ \cite{CarlsonMullerPeters}. We now construct a basis $\{\gamma_0, \gamma_1 \}$ for the homology group $H_1(\mathscr{E}_{\lambda},\mathbb{Z})$, which is isomorphic to $\mathbb{Z}^2$. Let $\gamma_0$ be a cycle on one copy of $\mathbb{C}$ that encircles the branch cut $(1,\infty)$; while let $\gamma_1$ be a circle that is the composite of two lines: the line from $1$ to $\lambda$ on the first copy of $\mathbb{C}$ and the line from $\lambda$ to $1$ on the second copy of $\mathbb{C}$. The integration of the oneform $\Omega_\lambda$ over the basis $\{\gamma_0, \gamma_1 \}$ gives us two periods $\varpi_0(\lambda)$ and $\varpi_1(\lambda)$
\begin{equation} \label{eq:Omega_lambdaIntegration}
\varpi_0(\lambda)= \int_{\gamma_0} \Omega_\lambda,~\varpi_1(\lambda)=\int_{\gamma_1} \Omega_\lambda,
\end{equation} 
which are multi-valued holomorphic functions. The computations of $\varpi_0(\lambda)$ and $\varpi_1(\lambda)$ are certainly very well-known and have a very long history. Up to a sign, they are given by the integrals
\begin{equation} \label{eq:periodsintegration}
\begin{aligned}
\varpi_0(\lambda)&=\frac{1}{\pi}\int_{1}^{\infty} \frac{dx}{\sqrt{x(x-1)(x-\lambda)}},\\
\varpi_1(\lambda)&=\frac{1}{\pi}\int_{1}^{\lambda}\, \frac{dx}{\sqrt{x(x-1)(x-\lambda)}}.
\end{aligned}
\end{equation}
After a change of variable by $x=1/z$, the first integral in formula \eqref{eq:periodsintegration} becomes
\begin{equation} \label{eq:periodZintegration}
\varpi_0(\lambda)=\frac{1}{\pi}\int_0^1 \frac{dz}{\sqrt{z(1-z)(1-\lambda z)}}.
\end{equation}
When $\lambda$ is in a small neighborhood of $0$, the factor $(1-\lambda z)^{-1/2}$ admits a power series expansion in $\lambda z$. Then the integration \eqref{eq:periodZintegration} becomes 
\begin{equation}
\varpi_0(\lambda)=\sum_{n=0}^\infty \binom{-1/2}{n}^2 \lambda^n,
\end{equation}
which is the hypergeometric function $_2F_1(1/2,1/2;1; \lambda)$ \cite{Chandra}. The period $\varpi_1(\lambda)$ can be computed similarly, and it admits an expansion with leading term given by
\begin{equation}
\varpi_1(\lambda)=-\frac{1}{\pi i} \int_{\lambda}^1 \frac{dx}{\sqrt{x(1-x)(x-\lambda)}}=-\frac{1}{\pi i}(4 \log 2- \log \lambda)+ \cdots,
\end{equation}
where the limit of the terms in $\cdots$ are zero when $\lambda \rightarrow 0$. Thus, we deduce that the period $\varpi_1(\lambda)$ must be of the form
\begin{equation}
\varpi_1(\lambda)=\frac{1}{\pi i} (\varpi_0(\lambda) \log \lambda+h(\lambda))-\frac{\log 16}{\pi i} \varpi_0(\lambda),
\end{equation}
where $h(\lambda)$ is a holomorphic function in a small neighborhood of $\lambda=0$ and it admits a power series expansion of the form \cite{CarlsonMullerPeters}
\begin{equation}
h(\lambda) =\frac{1 }{2}\,\lambda+\frac{21 }{64}\,\lambda^2+\frac{185 }{768}\,\lambda^3+ \cdots.
\end{equation}

The derivative of the oneform $\Omega_\lambda$ with respect to $\lambda$ is given by
\begin{equation}
\frac{d\,\Omega_\lambda}{d\lambda}=\frac{1}{4\pi} \frac{dx}{\sqrt{x(x-1)(x-\lambda)^3}},
\end{equation}
which is a meromorphic oneform on $\mathscr{E}_\lambda$ with a pole at the point $(x=\lambda,y=0)$. However, its residue at this pole is zero, hence it defines an element of $H^1(\mathscr{E}_{\lambda},\mathbb{C})$ through integration. Moreover, $\{\Omega_\lambda, \Omega'_\lambda \}$ form a basis for the two dimensional vector space $H^1(\mathscr{E}_{\lambda},\mathbb{C})$. The second derivative of $\Omega_\lambda$, i.e. $d^2\Omega/d\lambda^2$, also defines an element of $H^1(\mathscr{E}_{\lambda},\mathbb{C})$, hence it must be a linear combination of $\Omega_\lambda$ and $\Omega'_\lambda$. From the Chapter 1 of \cite{CarlsonMullerPeters}, the oneform $\Omega_\lambda$ satisfies a second order ordinary differential equation (ODE)
\begin{equation} \label{eq:picardfuchsequation}
\lambda(1-\lambda)\,\frac{d^2 \Omega}{d\lambda^2} +(1-2\lambda) \,\frac{d \,\Omega}{d \lambda}-\frac{1}{4} \,\Omega =0,
\end{equation}
which is called the Picard-Fuchs equation of $\Omega_\lambda$. Therefore, the periods $\varpi_i(\lambda), i=0,1$ also satisfy this ODE
\begin{equation} \label{eq:periodsPFequation}
\lambda(1-\lambda)\,\varpi''_i(\lambda) +(1-2\lambda)\, \varpi'_i(\lambda)-\frac{1}{4}\, \varpi_i(\lambda) =0, ~i=0,1.
\end{equation}
In fact, $\{\varpi_0(\lambda),\varpi_1(\lambda)\}$ form a basis for the solution space of the Picard-Fuchs equation \eqref{eq:picardfuchsequation}.

\section{ The modular lambda function and theta functions} \label{sec:modularlambdafunction}

In this section, we will review the properties of the modular lambda function, the periods $\varpi_i(\lambda), i=0,1$ and the theta functions. All the results in this section have been known since the 19th century, and the readers who are familiar with them can skip this section completely.

From the computations in Section \ref{sec:legendrefamily}, the monodromy of the periods $\{\varpi_0(\lambda), \varpi_1(\lambda)\}$ near the singular point $\lambda=0$ is
\begin{equation}
\varpi_0(\lambda) \rightarrow \varpi_0(\lambda), ~\varpi_1(\lambda) \rightarrow \varpi_1(\lambda)+2\, \varpi_0(\lambda).
\end{equation}
Let the column vector $\varpi$ be
\begin{equation}
\varpi=(\varpi_0,\varpi_1)^\top,
\end{equation}
then the monodromy matrix of $\varpi$ near  $\lambda=0$ is 
\begin{equation} \label{eq:monodromyaroundlambd0}
T_0=
\begin{pmatrix}
 1 & 0  \\
 2 & 1  \\
\end{pmatrix}.
\end{equation}
The monodromy matrices of the period vector $\varpi$ near the three singular points $\{0,1,\infty\}$ generate the modular group $\Gamma(2)$. The readers are referred to the book \cite{CarlsonMullerPeters} for more details. 

The period $\tau$ of the elliptic curve $\mathscr{E}_\lambda$ is given by the quotient
\begin{equation}\label{eq:canonicalTau}
\tau = \frac{\varpi_1(\lambda) }{\varpi_0(\lambda) }.
\end{equation}
In a small neighborhood of $\lambda=0$, $\tau$ has an explicit expansion of the form
\begin{equation} \label{eq:defnoftau}
\tau(\lambda)=\frac{1}{\pi i}\, \left(\log \lambda+\frac{h(\lambda) }{\varpi_0(\lambda) } \right)-\frac{\log 16}{\pi i}.
\end{equation}
The elliptic curve $\mathscr{E}_{\lambda}$ is isomorphic to the quotient of $\mathbb{C}$ by the lattice generated by $\{1, \tau \}$ \cite{CarlsonMullerPeters}. The inverse of $\tau(\lambda)$, i.e. $\lambda(\tau)$, is the famous modular lambda function. Moreover, $\lambda(\tau)$ generates the function field of the modular curve $X(2)$, i.e. it is a Hauptmodul for $X(2)$ \cite{Chandra}. The index of $\Gamma(2)$ in $\text{PSL}(2,\mathbb{Z})$ is 6 and the quotient group $\text{PSL}(2,\mathbb{Z})/\Gamma(2)$ is generated by 
\begin{equation}
T:\tau \mapsto \tau +1,~S:\tau \mapsto -\frac{1}{\tau}.
\end{equation}
The modular curve $X(2)=\Gamma(2)\backslash \mathbb{H}$ is a sixfold cover of $\text{PSL}(2,\mathbb{Z}) \backslash \mathbb{H}$ \cite{Diamond}.

Let the variable $q$ be defined by
\begin{equation}
q:=\exp \pi i \tau.
\end{equation}
From formula \eqref{eq:defnoftau}, we immediately obtain
\begin{equation} \label{eq:qintermsoflambda}
q=\frac{1}{16} \,\lambda\, \exp \left(h(\lambda)/\varpi_0(\lambda)\right).
\end{equation}
This equation can be inverted order by order, and the result is the power series expansion of $\lambda$ with respective to $q$ \cite{CarlsonMullerPeters}
\begin{equation} \label{eq:modularLambdafunction}
\lambda(q)=16 q-128 q^2 +704 q^3 -3072 q^4  + \cdots.
\end{equation}
The modular lambda function can also be expressed as a product of the theta functions \cite{Chandra}
\begin{equation} \label{eq:lambdaandtheta}
\lambda(q)=\frac{\theta_2^4(0,q)}{\theta_3^4(0,q)}, \,1-\lambda(q)=\frac{\theta_4^4(0,q)}{\theta_3^4(0,q)};
\end{equation}
where we have used the Jacobi identity
\begin{equation}
\theta_3^4(0,q)=\theta_2^4(0,q)+\theta_4^4(0,q).
\end{equation}

Following the mirror symmetry of Calabi-Yau threefolds \cite{MarkGross,KimYang}, let us define the normalized Yukawa coupling $\mathcal{Y}$ by
\begin{equation}
\mathcal{Y}:=\frac{1}{\varpi_0^2(\lambda) }\,\Big(\int_{\mathscr{E}_{\lambda}} \Omega \wedge \frac{d \Omega}{d \lambda}\Big) \,d \lambda,
\end{equation}
which is a oneform defined on $\mathbb{C}-\{0,1 \}$. In fact, the Yukawa coupling $\mathcal{Y}$ is just $d\tau$
\begin{equation} \label{eq:yukawacouplingdtau}
\mathcal{Y}=\frac{1}{\varpi_0^2(\lambda)}\,\left(\varpi_0(\lambda) \frac{d\varpi_1(\lambda)}{d\lambda}-\varpi_1(\lambda) \frac{d \varpi_0(\lambda)}{d \lambda} \right)\,d \lambda=\frac{d(\varpi_1(\lambda)/\varpi_0(\lambda))}{d\lambda} \,d\lambda=d \tau.
\end{equation}
Now define $W_k$ to be
\begin{equation}
W_k =\int_{\mathscr{E}_{\lambda}} \Omega \wedge \frac{d^k \Omega}{d \lambda^k}, ~k=0,1,2, \cdots.
\end{equation}
From the definition, we immediately have
\begin{equation}
W_0=0,~W_2 =\frac{dW_1}{d\lambda}.
\end{equation}
The Picard-Fuchs equation \eqref{eq:picardfuchsequation} implies
\begin{equation}
\lambda(\lambda-1) \frac{dW_1}{d\lambda}+(2 \lambda-1)\,W_1=0,
\end{equation}
a general solution of which is of the form
\begin{equation}
W_1 = \frac{C}{\lambda (1-\lambda)},
\end{equation}
Here $C$ is a nonzero constant, which can be shown to be $1/\pi i$ by an explicit computation using the series expansion of $\varpi_i(\lambda),~i=0,1$. Hence the normalized Yukawa coupling $\mathcal{Y}$ is also equal to
\begin{equation}
\mathcal{Y} = \frac{1}{\pi i }\, \frac{d \lambda}{\varpi_0^2(\lambda) \,\lambda(1-\lambda)}.
\end{equation}
Then formula \eqref{eq:yukawacouplingdtau} immediately implies
\begin{equation} \label{eq:piidtaudlambda}
\pi i \,\frac{d\tau}{d\lambda}=\frac{1}{\varpi_0^2(\lambda) \,  \lambda (1-\lambda) },
\end{equation}
which is equivalent to
\begin{equation} \label{eq:dlambdadtau}
\frac{1}{\pi i}\,\frac{d\lambda}{d \tau}=\varpi_0^2(\lambda)\,\lambda (1-\lambda).
\end{equation}

Let us now prove a well-know result \cite{Zagier, ZagierAT}.
\begin{lemma}
\begin{equation}
\varpi_0(\lambda(q))=\theta_3^2(0,q).
\end{equation}
\end{lemma}
\begin{proof}
From formula \eqref{eq:dlambdadtau}, we have
\begin{equation}\label{eq:varpi0squarelambda}
\varpi_0^2(\lambda)=\frac{1}{\pi i}\,\frac{d\lambda}{d \tau} \,\frac{1}{\lambda(1-\lambda)}.
\end{equation}
The actions of the two generators $T$ and $S$ of $\text{PSL}(2, \mathbb{Z})$ on $\lambda$ are given by
\begin{equation}
\begin{aligned}
T:&\,\lambda \mapsto \frac{\lambda}{\lambda-1}, \\
S:&\,\lambda \mapsto 1-\lambda.
\end{aligned}
\end{equation}
From formula \eqref{eq:varpi0squarelambda}, the actions of $T$ and $S$ on $\varpi_0^2(\lambda)$ are given by
\begin{equation}
\begin{aligned}
T:&\,\varpi_0^2(\lambda) \mapsto (1-\lambda)\,\varpi_0^2(\lambda), \\
S:&\,\varpi_0^2(\lambda) \mapsto -\tau^2\,\varpi_0^2(\lambda).
\end{aligned}
\end{equation}
But the actions of $T$ and $S$ on the $\theta^4_3(0,q)$ are \cite{Chandra}
\begin{equation}
\begin{aligned}
T:&\,\theta^4_3(0,q) \mapsto \theta^4_4(0,q)=(1-\lambda)\,\theta^4_3(0,q), \\
S:&\,\theta^4_3(0,q) \mapsto -\tau^2\,\theta^4_3(0,q);
\end{aligned}
\end{equation}
where we have used formula \eqref{eq:lambdaandtheta}. Since $\varpi_0(\lambda(q))$ does not vanish on the upper half plane, the quotient $\theta^4_3(0,q)/\varpi_0^2(\lambda(q))$ is a holomorphic function on $\mathbb{H}$ that is invariant under the actions of $\text{PSL}(2,\mathbb{Z})$. Near the cusp $i\infty$, the $q$-expansion of $\theta^2_3(0,q)/\varpi_0(\lambda(q))$ is given by
\begin{equation}
\theta^2_3(0,q)/\varpi_0(\lambda(q))=1+O(q^4),
\end{equation}
hence we deduce $\theta^2_3(0,q)/\varpi_0(\lambda(q))$ must be $1$.
\end{proof}

\noindent The upshot is that we have the following identities
\begin{equation} \label{eq:thetafnandlambdavarpi0}
\theta^4_2(0,q)=\lambda\,\varpi_0^2(\lambda(q)),\,\theta_3^2(0,q)=\varpi_0(\lambda(q)),\,\theta^4_4(0,q)=(1-\lambda)\,\varpi_0^2(\lambda(q)),
\end{equation}
which will be important in this paper.

\section{The meromorphic Seiberg-Witten oneform and its periods} \label{sec:seibergwittenperiods}

In this section, we will study the meromorphic Seiberg-Witten oneform $\Omega^{sw}_u$ \eqref{eq:swoneformU} and its periods. Under the transformation \eqref{eq:ellTransformation} which sends the family $\mathcal{E}_u$ \eqref{eq:swfamilyelliptic} to the Legendre family $\mathscr{E}_\lambda$ \eqref{eq:legendrefamilyequation} of elliptic curves, the meromorphic Seiberg-Witten oneform $\Omega^{sw}_u$ \eqref{eq:swoneformU} is sent to
\begin{equation} \label{eq:swoneformxdxovery}
\Omega^{sw}_\lambda=\frac{1}{\pi}\left( \frac{xdx}{y}-\frac{\lambda dx}{y} \right).
\end{equation}
Notice that $\Omega^{sw}_\lambda$ can also be expressed as
\begin{equation} \label{eq:swmeroomega}
\Omega^{sw}_\lambda=2\,(x \, \Omega_\lambda -\lambda \, \Omega_\lambda),
\end{equation}
where $\Omega_\lambda$ is the nowhere vanishing oneform \eqref{eq:LengendreOneform} on the elliptic curve $\mathscr{E}_\lambda$.

Let us now show that up to a total differential, $\Omega^{sw}_\lambda$ is equal to
\begin{equation} \label{eq:SWoneformLambda}
\Omega^{sw}_\lambda= 4\, \lambda (\lambda-1)\, \Omega'_\lambda + \frac{2}{ \pi} \,df,
\end{equation}
where $f$ is a meromorphic function on the elliptic curve $ \mathscr{E}_\lambda$
\begin{equation}
f:=x^{1/2}(x-1)^{1/2}(x-\lambda)^{-1/2}.
\end{equation}
The total differential of $f$ is
\begin{equation} 
\begin{aligned}
df=&\frac{1}{2}\, x^{-1/2}(x-1)^{1/2}(x-\lambda)^{-1/2}dx+\frac{1}{2} \, x^{1/2}(x-1)^{-1/2}(x-\lambda)^{-1/2}dx \\
      &-\frac{1}{2} \, x^{1/2}(x-1)^{1/2}(x-\lambda)^{-3/2}dx,
\end{aligned}
\end{equation}
whose residues at the singularities are all 0. The oneform $\Omega_\lambda$ and its derivative $\Omega'_\lambda$ are
\begin{equation}
\begin{aligned}
\Omega_\lambda&=\frac{1}{2 \pi} \, x^{-1/2}(x-1)^{-1/2}(x-\lambda)^{-1/2} \,dx, \\
\Omega'_\lambda&=\frac{1}{4 \pi } \,x^{-1/2}(x-1)^{-1/2}(x-\lambda)^{-3/2} \, dx, 
\end{aligned}
\end{equation}
from which we deduce
\begin{equation} \label{eq:dfomega}
\frac{1}{2 \pi} \,df=\frac{1}{2}\,(x-1)\, \Omega_\lambda + \frac{1}{2}\,x\, \Omega_\lambda-x\,(x-1)\, \Omega'_\lambda.
\end{equation}
To simplify this formula, we will need the identity
\begin{equation}
(x-\lambda)\, \Omega'_\lambda= \frac{1}{2} \,\Omega_\lambda,
\end{equation}
which tells us
\begin{equation}
x\,(x-1)\, \Omega'_\lambda=\frac{1}{2}\, x\, \Omega_\lambda +(\lambda-1)(\frac{1}{2} \,\Omega_\lambda + \lambda\, \Omega'_\lambda).
\end{equation}
Hence formula \eqref{eq:dfomega} becomes
\begin{equation}
\frac{1}{2 \pi} \,df=\frac{1}{2}\, x \,\Omega_\lambda-\frac{1}{2}\, \lambda\, \Omega_\lambda-\lambda(\lambda-1) \,\Omega'_\lambda,
\end{equation}
from which we immediately obtain the formula \eqref{eq:SWoneformLambda}.

The integrals of $\Omega^{sw}_\lambda$ with respect to the homology basis $\{ \gamma_0,\gamma_1 \}$ of $\mathscr{E}_\lambda$ constructed in Section \ref{sec:legendrefamily} are given by
\begin{equation} \label{eq:swperiodsdefinition}
\begin{aligned}
\pi_0(\lambda)&=\int_{\gamma_0}\,4 \,\lambda(\lambda-1) \,\Omega'_\lambda=4 \,\lambda(\lambda-1) \,\varpi'_0(\lambda), \\
\pi_1(\lambda)&=\int_{\gamma_1}\,4 \,\lambda(\lambda-1) \,\Omega'_\lambda= 4\,\lambda(\lambda-1) \,\varpi'_1(\lambda),
\end{aligned}
\end{equation}
where we have used formula \eqref{eq:Omega_lambdaIntegration}. From the Picard-Fuchs equation \eqref{eq:picardfuchsequation}, we deduce
\begin{equation} \label{eq:BPSperiodsDerivative}
(\pi_0)'=-\varpi_0,~(\pi_1)'=-\varpi_1.
\end{equation}
The periods $\pi_i(\lambda),~i=0,1$ will be called the BPS periods in this paper. Since the periods $\{a_D,a\}$ \eqref{eq:aaDperiods} are given by the integrals of the Seiberg-Witten oneform $\Omega^{sw}_u$ \eqref{eq:swoneformU} over an integral basis of $H_1(\mathcal{E}_u,\mathbb{Z})$, hence they are related to $\{\pi_0, \pi_1 \}$ by an integral linear transformation. In particular, the central charge lattice of the BPS states of the $N=2$ theory is also given by
\begin{equation} \label{eq:BPSmasslatticeLegendre}
\{m\,\pi_0+n\,\pi_1:(m,n) \in \mathbb{Z}^2 \}.
\end{equation}
After studying the behaviors of the limits of $\{\pi_0, \pi_1 \}$ near the singularities $\lambda=0,1, \infty$ and compare them with that of $\{a, a_D \}$ near $u=\pm 1, \infty$, which have been computed in \cite{SeibergWitten}, we find that \cite{SeibergWitten}
\begin{equation} \label{eq:fieldrelations}
a=-\pi_1-\pi_0,~ a_D=-\pi_1.
\end{equation}
Hence from formula \eqref{eq:CCCu}, the complexified coupling constant $\tau_{sw}$ of the $N=2$ theory is also given by
\begin{equation}
\tau_{sw} =\frac{\pi'_1}{\pi'_0+\pi'_1}=\frac{\varpi_1}{\varpi_0+\varpi_1}= \frac{\tau}{\tau + 1},
\end{equation}
where we have used formula \eqref{eq:BPSperiodsDerivative}.

The derivatives $\varpi'_i(\lambda), ~i=0,1$ are two linearly independent solutions for the hypergeometric differential equation
\begin{equation} \label{eq:BPS_PF}
\lambda (1-\lambda) \frac{d^2(\varpi'_i)}{d\lambda^2} +(2-4\lambda) \frac{d(\varpi'_i)}{d\lambda}-\frac{9}{4}\varpi'_i=0.
\end{equation}
In fact, $\varpi'_0(\lambda)$ is the hypergeometric function
\begin{equation}
\varpi'_0(\lambda) =\frac{1}{4} \,\, _2F_1(\frac{3}{2},\frac{3}{2};2; \lambda).
\end{equation}
On the other hand, the BPS periods $\pi_i(\lambda), ~i=0,1$ satisfy the following ODE
\begin{equation} \label{eq:qformPF}
\frac{d^2\pi_i}{d\lambda^2}+Q(\lambda) \pi_i =0,~i=0,1;
\end{equation}
where $Q(\lambda)$ is the rational expression
\begin{equation}
Q(\lambda) =-\frac{1}{4\lambda(1-\lambda)}.
\end{equation}
The ODE \ref{eq:qformPF} is the Q-form of the hypergeometric differential equation \ref{eq:BPS_PF}. This Q-form is closely related to Schwarzian derivative \cite{Hille}.

More explicitly, we will further define the Seiberg-Witten period $\varphi$ to be the quotient
\begin{equation} \label{eq:SeibergWittenPeriod}
\varphi(\lambda):=\pi_1(\lambda)/\pi_0(\lambda),
\end{equation}
which is also equal to
\begin{equation}
\varphi(\lambda) =\varpi'_1(\lambda)/\varpi'_0(\lambda).
\end{equation}
While the quotient $a_D/a$ can be expressed as
\begin{equation}
\frac{a_D}{a} =\frac{\pi_1}{\pi_0+\pi_1}=\frac{\varphi}{\varphi + 1}.
\end{equation}
There is a fundamental relation between the Schwarzian derivative and ODE in the complex plane, which immediately tells us that \cite{Hille}
\begin{equation} \label{eq:SDlambda}
\{\varphi, \lambda  \} =2 Q(\lambda).
\end{equation}
However, we will see that it is the Schwarzian derivative of $\varphi$ with respect to the variable $\tau$ of the upper half-plane $\mathbb{H}$ that is going to be more interesting and important in this paper \cite{SS}.

\section{ The Schwarzian derivative of the Seiberg-Witten period } \label{sec:SchwarzianDerivative}

In this section, we will first apply the results in Section \ref{sec:modularlambdafunction} to obtain the transformations of the BPS periods $\{ \pi_0, \pi_1 \}$ under the action of $\text{PSL}(2, \mathbb{Z})$. Then we will study the Schwarzian derivative of the Seiberg-Witten period $\varphi$ \eqref{eq:SeibergWittenPeriod} with respect to $\tau$, which is given by the theta function.

\subsection{The transformations of the BPS periods} \label{sec:BPSperiodsTransformation}

From formula \eqref{eq:canonicalTau}, the derivative of $\varpi_1$ with respect to $\lambda$ can also be written as
\begin{equation} \label{eq:varpi1prime}
\varpi'_1=(\tau \varpi_0)'=\frac{d\tau}{d\lambda}\,\varpi_0+\tau \,\varpi_0'=\frac{1}{\pi i} \frac{1}{\lambda(1-\lambda)\varpi_0} + \tau \,\varpi'_0,
\end{equation}
where we have used formula \eqref{eq:piidtaudlambda}. Under the action of $S: \tau \mapsto -1/\tau$, we have
\begin{equation}
S:\lambda \mapsto 1- \lambda;~d \lambda \mapsto -d\lambda;~\varpi_0 \mapsto i \tau \varpi_0=i \varpi_1;~\varpi_1=\tau\,\varpi_0 \mapsto -\frac{1}{\tau}\, i \tau \varpi_0 =-i\,\varpi_0;
\end{equation}
hence the transformations of $\{ \pi_0, \pi_1 \}$ under $S$ are given by
\begin{equation}
\begin{aligned}
S:\pi_0 &\mapsto -i \,\pi_1,\\
S:\pi_1 &\mapsto i\,\pi_0.
\end{aligned}
\end{equation}
Therefore the Seiberg-Witten period $\varphi$ \eqref{eq:SeibergWittenPeriod} transforms according to
\begin{equation}
S:\varphi \mapsto -\frac{1}{\varphi}.
\end{equation}

The transformation of $\{ \pi_0, \pi_1 \}$ under the action of $T$ is more complicated. Under the action of the transformation $T: \tau \mapsto \tau+1$, we have
\begin{equation}
T:\lambda \mapsto \frac{\lambda}{\lambda-1},~d\lambda \mapsto -\frac{1}{(\lambda-1)^2}\, d\lambda ,~\varpi_0 \mapsto \sqrt{1-\lambda} \,\varpi_0;
\end{equation}
from which we obtain
\begin{equation}
\begin{aligned}
T:\pi_0 & \mapsto\frac{2 \,\lambda }{\sqrt{1-\lambda } } \big(2 (\lambda -1) \varpi _0'+\varpi _0\big),\\
T:\pi_1  & \mapsto \frac{2 \, \lambda}{\sqrt{1-\lambda }}  \left(\left(2 (\lambda -1)  \frac{d\tau}{d\lambda} +\tau +1\right)\varpi _0+2 (\lambda -1) (\tau+1) \varpi _0'\right).
\end{aligned} 
\end{equation}
However under the action of $T^2: \tau \mapsto \tau+2$, we have 
\begin{equation}
T^2:\lambda \mapsto \lambda, ~d\lambda \mapsto d\lambda, ~\varpi_0 \mapsto \varpi_0,
\end{equation}
hence the BPS periods $\{ \pi_0, \pi_1 \}$  transform in the way
\begin{equation}
\begin{aligned}
 T^2:\pi_0 & \mapsto  \pi_0, \\
 T^2: \pi_1 & \mapsto  \pi_1+2 \,\pi_0.
 \end{aligned}
\end{equation}
Thus the Seiberg-Witten period $\varphi$ \eqref{eq:SeibergWittenPeriod} transforms in the way
\begin{equation}
T^2:\varphi \mapsto \varphi+2.
\end{equation}
The upshot is that $\varphi$ is equivariant with respect to the actions of $S$ and $T^2$.

\subsection{The fixed point of the $S$-transformation} \label{sec:fixedpoint}

The point $\lambda=1/2$, i.e. $u=0$, has an interesting property. Recall that the period $\tau$ of $\mathscr{E}_{\lambda}$ can also be expressed as \cite{Chandra}
\begin{equation}
\tau =i \frac{_2F_1(1/2,1/2;1; 1-\lambda)}{_2F_1(1/2,1/2;1; \lambda)},
\end{equation}
hence the value of $\tau$ at $\lambda = 1/2$ is
\begin{equation}
\tau|_{\lambda=1/2} = i.
\end{equation}
The elliptic curve $\mathscr{E}_{1/2}$ 
\begin{equation}
\mathscr{E}_{1/2}: y^2=x(x-1)(x-1/2)
\end{equation}
admits CM (complex multiplication), and its $j$-invariant is $1728$.

From the formula \eqref{eq:varpi1prime}, the Seiberg-Witten period $\varphi$ can also be expressed as
\begin{equation}
\varphi=\tau +\frac{1}{\pi i} \frac{1}{\lambda (1-\lambda) \varpi_0 \varpi'_0}.
\end{equation}
From the following special values of hypergeometric functions
\begin{equation}
\begin{aligned}
_2F_1(1/2,1/2;1; \lambda)|_{\lambda=1/2}&=\frac{\Gamma(1/4)}{(2 \pi)^{1/2} \Gamma(3/4)},\\
_2F_1(3/2,3/2;2; \lambda)|_{\lambda=1/2}&=-\frac{\Gamma(-1/4)}{(2 \pi)^{1/2} \Gamma(5/4)},
\end{aligned}
\end{equation}
we immediately obtain that the value of $\varphi$ at $\lambda =1/2$ is $-i$. Thus the central charge lattice of the BPS states \eqref{eq:BPSmasslatticeLegendre} has an additional symmetry. It is very interesting to find out whether this symmetry has any physical significance.

\subsection{The Schwarzian derivative}

Inspired by the paper \cite{SS}, let us now compute the Schwarzian derivative of the  Seiberg-Witten period $\varphi$ with respect to the variable $\tau$. From the results in Section \ref{sec:BPSperiodsTransformation},   $\varphi(\tau)$ is equivariant with respect to the modular group $\Gamma(2)$. First, $\varphi$ can be expressed as
\begin{equation} \label{eq:mirrormapnonperturbative}
\varphi =\frac{\varpi'_1}{\varpi'_0}=\frac{1}{\pi i} \left(\log \lambda+\frac{1}{\lambda} \frac{\varpi_0}{\varpi'_0}+\frac{h'}{\varpi'_0}-\log 16 \right).
\end{equation} 
An explicit computation of the first several terms of the $q$-expansion of $\{\varphi, \tau \}$, the transformation properties in the formula \ref{eq:SDvariabletransformation2} and results in the paper \cite{SS} show that 
\begin{equation}
\{\varphi, \tau \}=\frac{\pi^2}{2} \theta^8_3(0,q),~q=\exp(\pi i \tau).
\end{equation}
We have also proved this equation by using the property \eqref{eq:SDcocycle} of the Schwarzian derivative,  formula \eqref{eq:SDlambda} and the results in Section \ref{sec:modularlambdafunction}. Furthermore, using the properties \eqref{eq:SDMobiusTransformation} and \eqref{eq:SDvariabletransformation2}, we obtain
\begin{equation}
\{a_D/a, \tau_{sw} \}=\frac{\pi^2}{2} \theta^8_4(0,e^{\pi i \tau_{sw}}),
\end{equation}
where we have used the equation
\begin{equation}
 \theta^8_4 \left( 0,e^{\pi i \tau/(\tau+1)} \right) =(\tau +1)^4  \theta^8_3(0,e^{\pi i \tau}).
\end{equation}

The function $\theta^8_4(0,q)$ has an expansion of the form
\begin{equation}
\theta^8_4(0,q) = 1-16q+112q^2-448q^3+1136q^4-2016q^5+3136q^6+\cdots.
\end{equation}
On the other hand, the modular group $\Gamma_0(2)$ has a unique weight-4 normalized entire modular form $E_{0,4}(\tau)$ given by \cite{Diamond, OEIS}
\begin{equation}
E_{0,4}(\tau)=\eta^{16}(\tau)/\eta^8(2\tau).
\end{equation}
The modular form $E_{0,4}(\tau)$ is related to $\theta^8_4(0,q)$ by
\begin{equation}
\theta^8_4(0,q) =E_{0,4}(\tau/2).
\end{equation}
The readers are referred to the paper \cite{Borcherds} for some applications of the modular form $E_{0,4}(\tau)$.

\section{Great Picard's Theorem and the non-perturbative limit} \label{sec:greatpicards}

In this section, we will prove that the quotient $a_D/a$ does take real values in a small neighborhood of the nonperturbative limits $u=\pm 1$ using the Great Picard's Theorem. In fact, we will prove an equivalent result: the Seiberg-Witten period $\varphi$ \eqref{eq:SeibergWittenPeriod} takes real values near $\lambda = 0, 1 $.

First, let us look at the case where $\lambda=0$. Define $q^{sw}$ by
\begin{equation}
q^{sw}:=\exp \pi i\,\varphi.
\end{equation}
Thus $\varphi$ is real if and only if the modulus of $q^{sw}$ is 1. Formula \eqref{eq:mirrormapnonperturbative} implies
\begin{equation}
16\,q^{sw}=\lambda \exp (h'/\varpi'_0) \exp \left(\frac{1}{\lambda} \frac{\varpi_0}{\varpi'_0} \right).
\end{equation}
The power series expansion of $\varpi_0/\varpi'_0 $ in a small neighborhood of $\lambda=0$ is of the form
\begin{equation}
\varpi_0/\varpi'_0  =4-\frac{7}{2}\,\lambda+O(\lambda^2),
\end{equation}
from which we deduce that
\begin{equation}
16\,q^{sw}=F(\lambda) \exp (4/\lambda),~\text{with}~F(\lambda)=\lambda \exp (h'/\varpi'_0) \exp \left(\frac{1}{\lambda} \left( \frac{\varpi_0}{\varpi'_0}-4\right) \right).
\end{equation}
The function $F(\lambda)$ is holomorphic in a neighborhood of $\lambda=0$ and it admits a series expansion with the first term given by
\begin{equation}
F(\lambda)=\lambda+O(\lambda^2).
\end{equation}
The function $F(\lambda) \exp (4/\lambda)$ is holomorhpic in a small punctured neighborhood of $0$, while it has an essential singularity at $\lambda=0$. Now recall the following important theorem of Picard. 

\textbf{Great Picard's Theorem}: \textit{If $G(z)$ is a holomorphic function and $z_0$ is an essential singularity of $G(z)$,  then on any punctured neighborhood of $z_0$, $G$ takes on all possible complex values, with at most a single exception, infinitely many times.}

In particular, the \textbf{Great Picard's Theorem} implies that in every punctured neighborhood of $\lambda=0$, which is assumed to be small enough to avoid other singularities, the function $F(\lambda) \exp (4/\lambda)$ takes on all possible values in $\{\exp \pi i\,t: t \in \mathbb{R} \}$, with at most a single exception, infinitely many times. Therefore the behavior of $ \varphi$ in a small punctured neighborhood of $\lambda=0$ is of a very complicated nature. The behavior of the Seiberg-Witten period $\varphi$ at $\lambda=1$ can be analyzed similarly, which is related to the $\lambda=0$ case by an $S$-transformation.

\section{The scalar curvature of the moduli space} \label{sec:metric}

In this section, we will compute the scalar curvature of the moduli space $\mathscr{M}$ for the $N=2$ supersymmetric Yang-Mills theory. We will look at the behaviors of the scalar curvature near the singularities $u=\pm 1, \infty$.

From the formulas \eqref{eq:BPSperiodsDerivative} and \eqref{eq:fieldrelations}, the metric $(ds)^2$ \eqref{eq:metriconmodulispace} on the moduli space $\mathscr{M}$ can also be expressed as
\begin{equation} \label{eq:dsvarpi}
(ds)^2=-\frac{i}{2}(\varpi_1 \overline{\varpi}_0-\varpi_0\overline{\varpi}_1)d\lambda d\overline{\lambda}.
\end{equation}
We immediately recognize that this metric can be written down in a geometrically invariant way 
\begin{equation}
(ds)^2=(\frac{i}{2}\int_{\mathscr{E}_{\lambda}} \Omega_\lambda \wedge \overline{\Omega}_\lambda) d\lambda d\overline{\lambda}.
\end{equation}
The pull-back of the metric \eqref{eq:dsvarpi} to the upper half plane $\mathbb{H}$ becomes
\begin{equation}
(ds)^2=\left|\varpi_0 \frac{d\lambda}{d\tau} \right|^2~(\text{Im}\,\tau)d\tau d\overline{\tau},
\end{equation}
which is of a rather simple and explicit form. From the formula \eqref{eq:dlambdadtau}, we deduce that
\begin{equation} \label{eq:dstau}
(ds)^2=\pi^2 \left| \varpi_0^3 \lambda(1-\lambda) \right|^2 (\text{Im}\,\tau)d\tau d\overline{\tau}.
\end{equation}
The scalar curvature for the metric $(ds)^2$ \eqref{eq:dstau} is
\begin{equation}
R=\frac{1}{\pi^2\left| \varpi_0^3 \lambda(1-\lambda)  \right|^2 (\text{Im}\,\tau)^3 }.
\end{equation}
The scalar curvature $R$ is invariant under the action of $S$. While under the action of $T$, $R$ transforms according to
\begin{equation}
T:R \rightarrow \left| 1-\lambda \right|^{3} R. 
\end{equation}
Moreover, $R$ is invariant under the action of $T^2$. The scalar curvature $R$ has three singularities at $\lambda = 0, 1, \infty$. Let us now analyze the the behaviors of $R$ near them. When $\tau = i \infty$, we have $\lambda = 0$. On the imaginary axis, 
\begin{equation} \label{eq:imaginaryaxis}
\tau = i t,~t \in \mathbb{R}^+,
\end{equation}
when $t$ goes to infinity, the leading term of $R$ is
\begin{equation}
R \sim \frac{1}{256\pi^2 t^3  } \exp(2 \pi t),
\end{equation}
which approaches infinity exponentially. The behavior of $R$ near $\lambda=1$ ($\tau = 0$), is the same as that of $R$ near $\lambda=0$, which can be deduced from the property that $R$ is invariant under the $S$-transformation. Notice that the $S$-transformation maps a neighborhood of $\lambda = 0$ to a neighborhood of $\lambda=1$.

Now let us look at the behavior $R$ in a small neighborhood of $\lambda = \infty$. First do a $T$-transformation, which sends $\lambda= \infty$ to $\lambda =1$. Next do an $S$-transformation, which sends $\lambda=1$ to $\lambda=0$. Under the composition of these two transformations, $R$ transforms to
\begin{equation}
R \rightarrow |\lambda|^3 R=\frac{|\lambda|}{\pi^2\left| \varpi_0^3 (1-\lambda)  \right|^2 (\text{Im}\,\tau)^3}.
\end{equation}
On the imaginary axis \eqref{eq:imaginaryaxis}, when $t$ is large, the leading part of $|\lambda|^3R$ is 
\begin{equation}
|\lambda|^3 R \sim \frac{16}{\pi^2t^3} \exp(-\pi t),
\end{equation}
which approaches 0 exponentially. Hence we deduce that $R$ approaches 0 exponentially when $\lambda$ goes to infinity. So the moduli space $\mathscr{M}$ is asymptotically flat near $\lambda =\infty$, which reflects the property that $\lambda= \infty$ ($u=\infty$) is a perturbative limit of the $N=2$ supersymmetric Yang-Mills theory.

\section{Conclusions and further prospects } \label{sec:conclusion}

In this paper, we have applied the tools from number theory to the study of Seiberg-Witten theory. We have reviewed the properties of the periods of the Legendre family of elliptic curves, and their relations to the modular lambda function and the theta functions, which are included here in order to make this paper as self-contained as possible. Based on these results, we have studied the transformations of the Siberg-Witten period $\varphi$ \eqref{eq:SeibergWittenPeriod} under the action of the modular group $\text{PSL}(2, \mathbb{Z})$. We have shown that the Schwarzian derivative of $\varphi$ with respect to the period $\tau$ is given by the theta functions. This allows us to obtain the equation \eqref{eq:swperiodSchwarzian}, i.e. the Scharzian derivative of the quotient $a_D/a$ with respect to the complexified coupling constant $\tau_{sw}$ of the $N=2$ theory is also given by the theta functions.

We have also studied the behaviors of the quotient $a_D/a$ near the non-perturbative limit $u=\pm 1$ by using the Great Picard's Theorem. In particular, we have shown that $a_D/a$ can take on any real values, with at most a single exception, infinitely many times. The scalar curvature for the moduli space of the $N=2$ theory has also been computed. It has been shown that the scalar curvature approaches 0 exponentially near the perturbative limit $u=\infty$, so the moduli space is asymptotically flat near $u=\infty$. While the scalar curvature of the moduli space approaches infinity exponentially near the the two non-perturbative limit $u=\pm 1$.

There are several open questions left unanswered in this paper. Perhaps the most important open question is could the equation \eqref{eq:swperiodSchwarzian} bridge any new relations between the Seiberg-Witten theory and the theory of automorphic representations. For example, could we construct an automorphic representation from the quotient $a_D/a$ which is related to the counting of the BPS states of the $N=2$ supersymmetric Yang-Mills theory. It is also very interesting to see whether the equation \eqref{eq:swperiodSchwarzian} admits any physical interpretations.

As has been shown in Section \ref{sec:fixedpoint}, the point $u=0$ ($\lambda=1/2$) of the moduli space $\mathscr{M}$ is a special point. The lattice for the elliptic curve $\mathscr{E}_{1/2}$, which has $j$-invariant 1728, is 
\begin{equation}
\{m+n i:(m,n) \in \mathbb{Z}^2 \}.
\end{equation} 
While the central charge lattice of the $N=2$ theory for the vacuum with coordinate $u=0$ is isomorphic to this lattice. It is very interesting to see whether this property has any physical significance in the $N=2$ supersymmetric Yang-Mills theory.

\section*{Acknowledgments}

The author is grateful to Shamit Kachru and Arnav Tripathy for many helpful discussions and a reading of the draft.

\end{document}